Calculate the number of carbon atoms in a 90.0g sample of camphor C10H16O. Be sure your answer has a unit symbol if necessary, and round it to 3 significant digits.
Answer #1

Molar mass of C10H16O,

MM = 10*MM(C) + 16*MM(H) + 1*MM(O)

= 10*12.01 + 16*1.008 + 1*16.0

= 152.228 g/mol

mass(C10H16O)= 90.0 g

use:

number of mol of C10H16O,

n = mass of C10H16O/molar mass of C10H16O

=(90 g)/(1.522*10^2 g/mol)

= 0.5912 mol

use:

number of molecules = number of mol * Avogadro’s number

number of molecules = 0.5912 * 6.022*10^23 molecules

number of molecules = 3.56*10^23 molecules

1 C10H16O has 10 atoms of C

So,

Number of C = 10 * number of molecules

= 10 * 3.56*10^23 C atoms

= 3.56*10^24 C atoms

Answer: 3.56*10^24